# Invariant Random Forest: Tree-Based Model Solution for OOD Generalization

### 07 Dec 2023 v1

@@ 1 Introduction
The performance of existing machine learning models may drop severely in OOD scenarios (Shen et al. 2021; Zhou et al. 2021; Geirhos et al. 2020). To make up for this weakness, many methods have been proposed to improve the OOD generalization ability of machine learning models, such as IRM (Arjovsky et al. 2019) and REx (Krueger et al. 2021). But almost all the methods are caring about Deep Neural Networks (DNNs). For other types of models in machine learning, there are no solutions proposed so far.
@@ 4.1 Invariant Classification Tree
where $Q_{m}^{e}$ is the training data set from the $e$-th environment, $L$ is the penalty term regarding the invariance among environments, and $\lambda\in\mathbb{R}_{+}$ is the penalty weight. Here, we define $L$ to be the ratio between the highest and the lowest $I(Q_{m}^{e},\theta)$, which is
$L(Q_{m}^{1},...,Q_{m}^{E},\theta)=\max\limits_{e,f}I(Q_{m}^{e},\theta)/I(Q_{m}^{f},\theta).$ (16)
@@ 5 Experiments
Table 1: Accuracy and cross-entropy loss results of the synthetic classification task. We consider four cases when the dataset dimension varies in $\{2,5,10,20\}$. The numbers reported are the averages from 5 trials with different randomness seeds.
Dataset Dimension RF XGBoost IRF ($\lambda=1$) IRF ($\lambda=5$) IRF ($\lambda=10$)
$d=2$ 48.74 (0.83) 49.50 (1.24) 50.24 (0.75) 51.20 (0.73) 51.06 (0.73)
$d=5$ 47.62 (0.85) 49.14 (1.49) 52.24 (0.74) 55.04 (0.70) 55.12 (0.70)
$d=10$ 43.26 (0.86) 46.64 (1.68) 51.26 (0.74) 53.06 (0.71) 54.94 (0.70)
$d=20$ 40.08 (0.84) 48.28 (1.51) 52.56 (0.71) 55.08 (0.69) 57.42 (0.68)
@@
Dataset Dimension RF XGBoost IRF ($\lambda=1$) IRF ($\lambda=5$) IRF ($\lambda=10$)
$d=2$ 1.000 1.190 (0.033) 0.940 (0.017) 0.893 (0.019) 0.889 (0.024)
$d=5$ 1.000 1.041 (0.035) 0.874 (0.017) 0.826 (0.022) 0.847 (0.015)
$d=10$ 1.000 1.054 (0.034) 1.049 (0.049) 0.881 (0.013) 0.900 (0.021)
$d=20$ 1.000 1.052 (0.023) 1.042 (0.022) 0.916 (0.023) 0.934 (0.021)
Next, we test the proposed Invariant Random Forest (IRF) on several datasets with different OOD settings. We compare our method with Random Forest (RF) and XGBoost, which are two most popular ensemble tree models that are even being used in industries very often. Other methods are not included for comparisons because there are few works considering the OOD generalization of tree models, as far as we know.
If not stated otherwise, for results of classification tasks, the first number in the table represents the average accuracy in percentage, and the number in the bracket is the average cross-entropy loss. For results of regression tasks, the first number represents the average Mean Square Error (MSE), and the number in the bracket is the standard deviation in multiple runs. For regression tasks, we standardize the MSE results of RF to 1, and the MSE of every other method is shown using the ratio compared to RF. That is to say, a number in the table of a regression task less than 1 means better performance than RF.
@@ 5.1 Synthetic Data Experiments
$Y$ $=1_{d}^{\top}X_{1}+\mathcal{N}(0,1),$
$X_{2}$ ={$&Y,\qquad\quad\text{if }e=1,\\ &\mathcal{N}(0,1),\quad\text{otherwise}.\\ \end{aligned}\right.$
Again the training data is $e=1,2$, and the testing data is the environment $e=3$. In environment $e=1$, $X_{2}$ can predict $Y$ exactly without any error. But in other environments, $X_{2}$ has no predictability to $Y$ at all. Models that use $X_{1}$ to predict $Y$ would achieve greater performance in the testing environment, because it is more stable.
The results are shown in Table 1 and 2. In both tasks, IRF shows superior performance compared to the other methods. This proves that our method can recognize the unstable splits and avoid using these splits during the growing of the trees. A deep looking into the details reveals that our method uses less of $X_{2}$ and more of $X_{1}$ to predict than the others, as shown in the appendix.
@@ 5.2 Settings of Real Data Experiments
Table 3: Accuracy and cross-entropy loss results of real-data classification tasks in Scenario 1.
Dataset RF XGBoost IRF ($\lambda=1$) IRF ($\lambda=5$) IRF ($\lambda=10$)
Financial 53.07 (0.73) 53.05 (0.98) 55.30 (0.70) 56.34 (0.69) 56.84 (0.69)
Technical 49.07 (0.72) 49.10 (1.10) 49.36 (0.72) 50.07 (0.72) 50.65 (0.72)
Dataset RF XGBoost IRF ($\lambda$ = 1) IRF ($\lambda$ = 5) IRF ($\lambda$ = 10)
Energy1-1 1.000 1.266 (0.153) 1.009 (0.007) 1.029 (0.016) 1.027 (0.024)
Energy1-2 1.000 1.201 (0.102) 0.986 (0.027) 0.961 (0.086) 0.958 (0.109)
Energy1-3 1.000 1.219 (0.106) 1.018 (0.012) 1.032 (0.004) 1.045 (0.023)
Energy1-4 1.000 1.287 (0.125) 0.933 (0.084) 0.896 (0.129) 0.892 (0.129)
Energy2-1 1.000 1.186 (0.095) 0.891 (0.131) 0.862 (0.206) 0.854 (0.218)
Energy2-2 1.000 1.088 (0.022) 0.997 (0.017) 0.976 (0.083) 0.974 (0.094)
Energy2-3 1.000 1.144 (0.103) 0.973 (0.015) 1.019 (0.075) 1.054 (0.091)
Energy2-4 1.000 1.010 (0.063) 0.988 (0.019) 1.003 (0.007) 0.983 (0.039)
Air1-1 1.000 1.130 (0.015) 0.878 (0.122) 0.850 (0.200) 0.865 (0.222)
Air1-2 1.000 1.088 (0.043) 0.969 (0.016) 0.911 (0.025) 0.886 (0.063)
Air1-3 1.000 1.107 (0.076) 0.983 (0.027) 0.944 (0.085) 0.923 (0.084)
Air1-4 1.000 1.243 (0.061) 0.901 (0.148) 0.974 (0.228) 1.006 (0.246)
Air2-1 1.000 1.214 (0.084) 0.987 (0.003) 0.976 (0.024) 0.995 (0.043)
Air2-2 1.000 1.260 (0.122) 0.981 (0.033) 1.013 (0.064) 1.092 (0.068)
Air2-3 1.000 1.242 (0.144) 0.973 (0.020) 0.965 (0.035) 0.984 (0.026)
Air2-4 1.000 0.858 (0.020) 1.037 (0.009) 1.076 (0.012) 1.057 (0.010)
The ways of obtaining initial environment partitions and the settings of training and testing sets are different for classification and regression. We introduce them separately in the following subsections. The average results of IRF from multiple runs of training and testing are reported and are compared to RF and XGBoost. In Scenario 1, there is no hyper-parameter tuning and we set $\lambda=1,5,10$, respectively. For the validation procedure in Scenario 2 and 3, we first choose the proper maximum tree depth for RF on validation set. Then, we use the same maximum depth in IRF and use the validation set to choose the best $\lambda$ for IRF. For details regarding the validation procedure and the choices of some other hyper-parameters, please go to the appendix.
5.3 Real-Data Classification Tasks
Financial
Technical
Asset Pricing
Table 5: Accuracy and cross-entropy loss results of real-data classification tasks in Scenario 2 and 3. S2 is the upper part and S3 is the lower part.
Dataset RF XGBoost IRF
Financial 54.42 (0.71) 52.46 (0.85) 56.20 (0.69)
Technical 48.89 (0.71) 49.05 (1.14) 49.18 (0.72)
Asset Pricing 52.50 (0.70) 51.95 (0.75) 52.92 (0.69)
Financial 54.50 (0.70) 52.46 (0.85) 54.27 (0.71)
Technical 48.94 (0.72) 49.05 (1.14) 49.57 (0.72)
Asset Pricing 52.64 (0.70) 51.95 (0.75) 52.79 (0.69)
Dataset RF XGBoost IRF
Energy1-1 1.000 1.064 (0.113) 0.989 (0.047)
Energy1-2 1.000 1.141 (0.231) 0.925 (0.116)
Energy2-1 1.000 1.027 (0.107) 1.006 (0.050)
Energy2-2 1.000 1.059 (0.103) 1.037 (0.121)
Air1-1 1.000 1.052 (0.099) 0.916 (0.106)
Air1-2 1.000 1.112 (0.097) 0.979 (0.064)
Air1-3 1.000 1.173 (0.191) 0.940 (0.141)
Air1-4 1.000 1.099 (0.042) 1.006 (0.013)
Air2-1 1.000 1.187 (0.139) 0.970 (0.078)
Air2-2 1.000 1.191 (0.117) 0.982 (0.062)
Air2-3 1.000 1.236 (0.109) 0.986 (0.042)
Air2-4 1.000 1.089 (0.084) 0.995 (0.031)
Energy1-1 1.000 1.050 (0.086) 0.977 (0.056)
Energy1-2 1.000 1.187 (0.200) 0.979 (0.142)
Energy2-1 1.000 1.008 (0.097) 1.038 (0.143)
Energy2-2 1.000 1.068 (0.099) 1.062 (0.174)
Air1-1 1.000 1.053 (0.076) 0.845 (0.230)
Air1-2 1.000 1.126 (0.076) 0.956 (0.101)
Air1-3 1.000 1.151 (0.202) 0.988 (0.072)
Air1-4 1.000 1.110 (0.046) 0.993 (0.010)
Air2-1 1.000 1.173 (0.115) 0.926 (0.180)
Air2-2 1.000 1.205 (0.126) 1.005 (0.070)
Air2-3 1.000 1.223 (0.108) 1.005 (0.055)
Air2-4 1.000 1.091 (0.076) 0.991 (0.031)
5.4 Real-Data Regression Tasks
The training and testing settings of regression tasks all follow a same strategy. For the first scenario, we obtain three environments using the variable containing meta information first. We further use any two environments as training set, and the other one as testing set (3 runs for each dataset). For the second and third scenarios, we obtain four environments initially and use any two as training set, any other one as validation set, and the remaining one as testing set (12 runs for each dataset). We introduce the four datasets used as follow.
Energy1 & Energy2
Energy datasets [^3] https://github.com/LuisM78/Appliances-energy-prediction-data/blob/master/energydata_complete.csv, https://www.kaggle.com/datasets/gmkeshav/tetuan-city-power-consumption provide some information like temperature, humidity, and windspeed in a city or a room. We want to predict the energy consumption of that location. Since time has a heterogeneous influence on the matter of energy consumption (e.g., a big difference between day and night), in both datasets, we use the variable hour to split the data and obtain environments. There are four different splitting strategies considered for Scenario 1 and two strategies for Scenario 2 and 3, resulting in more derived datasets named as Energy1-$i$ (or Energy2-$i$). The details of the splitting strategies are in the appendix.
Air1 & Air2
PRSA_data_2010.1.1-2014.12.31.csv,
contain various types of air quality data, and models need to predict the PM2.5 index based on the air quality data given. In these two datasets, we use the variable month to split and obtain environments, because in different seasons, the behaviors of air pollution may be different. We again have different splitting strategies considered (four for Scenario 1, 2, and 3), resulting in more derived datasets named as Air1-$i$ (or Air2-$i$).
Results
6 Conclusion
Appendix A Experiment Details
A.1 Dataset Preprocessing
A.2 Environment Partitions for Regression Datasets
For regression datasets in real data experiments, we first split the whole dataset into 3 or 4 environments by splitting a certain variable. Training set, validation set (if needed), and testing set take the data from different environments, so these sets have heterogeneous distributions. In this way, the performance on the testing set can reflect the OOD generalization ability of different models. The construction of these sets is described in the main content of the paper. However, it has more than one strategy to split the certain variable, resulting in more than one derived datasets. Table 7 lists the details of how we split each dataset into multiple environments.
Table 7: Splitting strategies of regression datasets. Each dataset is split into three or four environments by the strategies in the table. The training set, validation set, and testing set are constructed by the rule described in the main content of the paper.
@@ A.2 Environment Partitions for Regression Datasets
A.3 Training Procedure
Appendix B More Analysis on Synthetic Data Experiments
Stable Environmental Stable Environmental Stable Environmental Stable Environmental
Regression ($d=2$) 7.58 (0.06) 4.05 (0.05) 9.15 (0.13) 2.11 (0.11) 9.89 (0.06) 1.42(0.03) 10.01 (0.05) 1.38 (0.03)
Regression ($d=5$) 8.20 (0.08) 4.00 (0.04) 9.80 (0.13) 1.51 (0.21) 10.47 (0.02) 0.73 (0.01) 10.56 (0.04) 0.69 (0.03)
Regression ($d=10$) 8.27 (0.10) 3.87 (0.04) 8.89 (0.39) 2.85 (0.54) 10.80 (0.06) 0.52 (0.07) 10.90 (0.06) 0.41 (0.02)
Regression ($d=20$) 8.45 (0.04) 3.81 (0.06) 9.16 (0.22) 2.46(0.35) 11.09 (0.04) 0.28 (0.03) 11.16 (0.05) 0.23 (0.01)

### 20 Dec 2023 v2

@@ 1 Introduction
The performance of existing machine learning models may drop severely in OOD scenarios (Shen et al. 2021; Zhou et al. 2022; Geirhos et al. 2020). To make up for this weakness, many methods have been proposed to improve the OOD generalization ability of machine learning models, such as IRM (Arjovsky et al. 2019) and REx (Krueger et al. 2021). But almost all the methods are caring about Deep Neural Networks (DNNs). For other types of models in machine learning, there are no solutions proposed so far.
@@ 4.1 Invariant Classification Tree
where $Q_{m}^{e}$ is the training data set from the $e$-th environment, $L$ is the penalty term regarding the invariance among environments, and $\lambda\in\mathbb{R}_{+}$ is the penalty weight. Notice that if $L\geq 0$, then $\lambda\rightarrow\infty$ will push $L\rightarrow 0$. Here, we define $L$ to be the ratio between the highest and the lowest $I(Q_{m}^{e},\theta)$ minus 1, which is
$L(Q_{m}^{1},...,Q_{m}^{E},\theta)=\max\limits_{e,f}I(Q_{m}^{e},\theta)/I(Q_{m}^{f},\theta)-1.$ (16)
As a result, when $\lambda$ becomes larger, the difference among $I(Q_{m}^{e},\theta)$ from different environments should be smaller when minimizing the objective function (15). When $\lambda\rightarrow\infty$, $L$ is pushed to be 0, meaning that all $I(Q_{m}^{e},\theta)$ will be the same, which is exactly the hard restriction in (14).
@@ 5 Experiments
Table 1: Accuracy results of the synthetic classification task. We consider four cases when the dataset dimension varies in $\{2,5,10,20\}$. The numbers reported are the averages from 5 trials with different randomness seeds. Standard deviations are in the brackets.
Dataset Dimension RF LR IRM REx XGBoost IRF ($\lambda=1$) IRF ($\lambda=5$) IRF ($\lambda=10$)
$d=2$ 48.74 (0.92) 47.22 (0.80) 44.90 (2.92) 52.72 (1.24) 49.50 (1.22) 50.24 (0.96) 51.20 (0.65) 51.06 (0.62)
$d=5$ 47.62 (0.61) 48.62 (0.24) 47.54 (0.68) 53.54 (0.46) 49.14 (0.41) 52.24 (0.86) 55.04 (0.82) 55.12 (1.22)
$d=10$ 43.26 (0.90) 47.12 (0.40) 47.86 (1.89) 53.24 (0.94) 46.64 (0.42) 51.26 (1.33) 53.06 (0.89) 54.94 (0.58)
$d=20$ 40.08 (0.79) 51.78 (0.68) 47.44 (1.01) 54.56 (0.89) 48.28 (0.62) 52.56 (1.02) 55.08 (0.75) 57.42 (0.47)
@@
Dataset Dimension RF LR IRM REx XGBoost IRF ($\lambda=1$) IRF ($\lambda=5$) IRF ($\lambda=10$)
$d=2$ 1.0 2.095 (0.056) 1.014 (0.027) 1.076 (0.062) 1.131 (0.020) 0.982 (0.014) 0.914 (0.020) 0.796 (0.014)
$d=5$ 1.0 1.779 (0.016) 0.786 (0.031) 0.928 (0.018) 1.028 (0.019) 1.015 (0.010) 0.956 (0.019) 0.761 (0.022)
$d=10$ 1.0 1.486 (0.030) 0.864 (0.033) 0.807 (0.013) 1.008 (0.042) 1.022 (0.009) 0.930 (0.014) 0.721 (0.016)
$d=20$ 1.0 1.348 (0.029) 0.895 (0.022) 0.818 (0.015) 1.066 (0.018) 0.990 (0.009) 0.874 (0.008) 0.684 (0.011)
Stable Environ- mental Stable Environ- mental Stable Environ- mental Stable Environ- mental
Regression ($d=2$) 6.54 (0.03) 4.87 (0.02) 6.69 (0.18) 4.65 (0.16) 8.15 (0.18) 3.56(0.22) 9.26 (0.08) 2.17 (0.11)
Regression ($d=5$) 7.52 (0.02) 4.01 (0.02) 7.43 (0.07) 3.91 (0.07) 8.75 (0.09) 3.00 (0.06) 9.72 (0.06) 1.75 (0.06)
Regression ($d=10$) 7.95 (0.02) 3.65 (0.01) 7.88 (0.05) 3.52 (0.04) 9.12 (0.07) 2.54 (0.10) 10.13 (0.06) 1.34 (0.07)
Regression ($d=20$) 8.31 (0.06) 3.42 (0.04) 8.21 (0.04) 3.27 (0.03) 9.54 (0.06) 2.16 (0.07) 10.49 (0.02) 1.03 (0.02)
Table 4: Accuracy results of real-data classification tasks in Scenario 1. Standard deviations are in the brackets.
Dataset RF XGBoost IRF ($\lambda=1$) IRF ($\lambda=5$) IRF ($\lambda=10$)
Financial 53.07 (1.06) 53.05 (1.27) 55.30 (1.27) 56.34 (1.42) 56.84 (1.47)
Technical 49.07 (0.29) 49.10 (0.33) 49.36 (0.45) 50.07 (0.67) 50.65 (0.75)
Table 5: Accuracy results of real-data classification tasks in Scenario 2 and 3. S2 is the upper part and S3 is the lower part. Standard deviations are in the brackets.
Dataset RF XGBoost IRF
Financial 54.42 (1.19) 52.46 (1.32) 56.20 (1.44)
Technical 48.89 (0.71) 49.05 (0.72) 49.18 (1.14)
Asset Pricing 52.50 (0.50) 51.95 (0.33) 52.92 (0.47)
Financial 54.50 (1.33) 52.46 (1.32) 54.27 (1.22)
Technical 48.94 (1.79) 49.05 (1.69) 49.57 (2.06)
Asset Pricing 52.64 (0.49) 51.95 (0.33) 52.79 (0.44)
Next, we test the proposed Invariant Random Forest (IRF) on several datasets with different OOD settings. We compare our method with Random Forest (RF) and XGBoost, which are two most popular ensemble tree models that are even being used in industries very often. For synthetic data experiments, we also include the results of linear regression (LR) and two popular OOD generalization methods on neural networks, IRM and REx, for reference. Other methods are not included for comparisons because there are few works considering the OOD generalization of tree models, as far as we know.
If not stated otherwise, for results of classification tasks, the first number in the table represents the average accuracy in percentage, and the number in the bracket is the standard deviation in multiple runs. For results of regression tasks, the first number represents the average Mean Square Error (MSE), and the number in the bracket is the standard deviation in multiple runs. For regression tasks, we standardize the MSE results of RF to 1, and the MSE of every other method is shown using the ratio compared to RF. That is to say, a number in the table of a regression task less than 1 means better performance than RF.
5.1 Hyperparameter Settings
@@ 5.2 Synthetic Data Experiments
$Y$ $=1_{d}^{\top}X_{1}+N,\quad N\sim\mathcal{N}(0,d),$
$X_{2}$ $=Y+N_{e},\quad N_{e}\sim\mathcal{N}(0,\sigma_{e}^{2}d).$
Again the training data is $e=1,2$, and the testing data is the environment $e=3$. We set $\sigma_{1}=0.1$, $\sigma_{2}=2$, and $\sigma_{3}=5$. In environment $e=1$, $X_{2}$ can predict $Y$ directly with a relatively low error. But in other environments, $X_{2}$ has high errors in predicting $Y$. Models that use $X_{1}$ to predict $Y$ would achieve greater performance in the testing environment, because it is more stable.
The results are shown in Table 1 and 2. In both tasks, IRF shows superior performance compared to the other methods, across different dataset dimensions. This proves that our method can recognize the unstable splits and avoid using these splits during the growing of the trees. Moreover, a large $\lambda$ seems to be a good choice. In real applications, $\lambda$ can be selected with a validation set. Actually, when $\lambda$ increases, our method uses less of $X_{2}$ and more of $X_{1}$ to predict, as illustrated as follows.
@@ 5.3 Settings of Real Data Experiments
Table 6: Splitting strategies of regression datasets. Each dataset is split into three or four environments by the strategies in the table. The training set, validation set, and testing set are constructed using these environments.
@@
Dataset RF XGBoost IRF ($\lambda$ = 1) IRF ($\lambda$ = 5) IRF ($\lambda$ = 10)
Energy1-1 1.000 1.266 (0.153) 1.009 (0.007) 1.029 (0.016) 1.027 (0.024)
Energy1-2 1.000 1.201 (0.102) 0.986 (0.027) 0.961 (0.086) 0.958 (0.109)
Energy1-3 1.000 1.219 (0.106) 1.018 (0.012) 1.032 (0.004) 1.045 (0.023)
Energy1-4 1.000 1.287 (0.125) 0.933 (0.084) 0.896 (0.129) 0.892 (0.129)
Energy2-1 1.000 1.186 (0.095) 0.891 (0.131) 0.862 (0.206) 0.854 (0.218)
Energy2-2 1.000 1.088 (0.022) 0.997 (0.017) 0.976 (0.083) 0.974 (0.094)
Energy2-3 1.000 1.144 (0.103) 0.973 (0.015) 1.019 (0.075) 1.054 (0.091)
Energy2-4 1.000 1.010 (0.063) 0.988 (0.019) 1.003 (0.007) 0.983 (0.039)
Air1-1 1.000 1.130 (0.015) 0.878 (0.122) 0.850 (0.200) 0.865 (0.222)
Air1-2 1.000 1.088 (0.043) 0.969 (0.016) 0.911 (0.025) 0.886 (0.063)
Air1-3 1.000 1.107 (0.076) 0.983 (0.027) 0.944 (0.085) 0.923 (0.084)
Air1-4 1.000 1.243 (0.061) 0.901 (0.148) 0.974 (0.228) 1.006 (0.246)
Air2-1 1.000 1.214 (0.084) 0.987 (0.003) 0.976 (0.024) 0.995 (0.043)
Air2-2 1.000 1.260 (0.122) 0.981 (0.033) 1.013 (0.064) 1.092 (0.068)
Air2-3 1.000 1.242 (0.144) 0.973 (0.020) 0.965 (0.035) 0.984 (0.026)
Air2-4 1.000 0.858 (0.020) 1.037 (0.009) 1.076 (0.012) 1.057 (0.010)
Dataset RF XGBoost IRF
Energy1-1 1.000 1.064 (0.113) 0.989 (0.047)
Energy1-2 1.000 1.141 (0.231) 0.925 (0.116)
Energy2-1 1.000 1.027 (0.107) 1.006 (0.050)
Energy2-2 1.000 1.059 (0.103) 1.037 (0.121)
Air1-1 1.000 1.052 (0.099) 0.916 (0.106)
Air1-2 1.000 1.112 (0.097) 0.979 (0.064)
Air1-3 1.000 1.173 (0.191) 0.940 (0.141)
Air1-4 1.000 1.099 (0.042) 1.006 (0.013)
Air2-1 1.000 1.187 (0.139) 0.970 (0.078)
Air2-2 1.000 1.191 (0.117) 0.982 (0.062)
Air2-3 1.000 1.236 (0.109) 0.986 (0.042)
Air2-4 1.000 1.089 (0.084) 0.995 (0.031)
Energy1-1 1.000 1.050 (0.086) 0.977 (0.056)
Energy1-2 1.000 1.187 (0.200) 0.979 (0.142)
Energy2-1 1.000 1.008 (0.097) 1.038 (0.143)
Energy2-2 1.000 1.068 (0.099) 1.062 (0.174)
Air1-1 1.000 1.053 (0.076) 0.845 (0.230)
Air1-2 1.000 1.126 (0.076) 0.956 (0.101)
Air1-3 1.000 1.151 (0.202) 0.988 (0.072)
Air1-4 1.000 1.110 (0.046) 0.993 (0.010)
Air2-1 1.000 1.173 (0.115) 0.926 (0.180)
Air2-2 1.000 1.205 (0.126) 1.005 (0.070)
Air2-3 1.000 1.223 (0.108) 1.005 (0.055)
Air2-4 1.000 1.091 (0.076) 0.991 (0.031)
The ways of obtaining initial environment partitions and the settings of training and testing sets are different for classification and regression. We introduce them separately in the following subsections. The average results of IRF from multiple runs of training and testing are reported and are compared to RF and XGBoost. In Scenario 1, there is no hyper-parameter tuning and we set $\lambda=1,5,10$, respectively. For the validation procedure in Scenario 2 and 3, we first choose the proper maximum tree depth for RF on validation set. Then, we use the same maximum depth in IRF and use the validation set to choose the best $\lambda$ for IRF. For details regarding the validation procedure and the choices of some other hyper-parameters, please go to Section 5.1.
5.4 Real-Data Classification Tasks
Financial
Technical
Asset Pricing
5.5 Real-Data Regression Tasks
The training and testing settings of regression tasks all follow a same strategy. For the first scenario, we obtain three environments using the variable containing meta information first. We further use any two environments as training set, and the other one as testing set (3 runs for each dataset). For the second and third scenarios, we obtain four environments initially and use any two as training set, any other one as validation set, and the remaining one as testing set (12 runs for each dataset). In this way, the training set and testing set will have heterogeneous distributions, hence the performance on the testing set can reflect the OOD generalization ability of the model. We introduce the four datasets used as follow.
Energy1 & Energy2
Energy datasets [^3] https://github.com/LuisM78/Appliances-energy-prediction-data/blob/master/energydata_complete.csv, https://www.kaggle.com/datasets/gmkeshav/tetuan-city-power-consumption provide some information like temperature, humidity, and windspeed in a city or a room. We want to predict the energy consumption of that location. Since time has a heterogeneous influence on the matter of energy consumption (e.g., a big difference between day and night), in both datasets, we use the variable hour to split the data and obtain environments. There are four different splitting strategies considered for Scenario 1 and two strategies for Scenario 2 and 3, resulting in more derived datasets named as Energy1-$i$ (or Energy2-$i$). The details of the splitting strategies are in Table 6.
Air1 & Air2
PRSA_data_2010.1.1-2014.12.31.csv,
contain various types of air quality data, and models need to predict the PM2.5 index based on the air quality data given. In these two datasets, we use the variable month to split and obtain environments, because in different seasons, the behaviors of air pollution may be different. We again have different splitting strategies considered (four for Scenario 1, 2, and 3), resulting in more derived datasets named as Air1-$i$ (or Air2-$i$), as shown in Table 6.
Results
6 Conclusion